OpenStudy (mymakersdaughter):

How do I find the roots and multiplicity of this: y=4x^3-2x^2-2x? I know that the answer is 0 and -2 with a multiplicity of 3 but i need to know how to do it

hartnn (hartnn):

start by factoring the 'x' from all terms... you'll get an easy quadratic as another factor :)

OpenStudy (mymakersdaughter):

so that would be x(4x^2-2x-2)?

hartnn (hartnn):

yep, notice you could factor out the 2 as well. \(2x(2x^2 -x-1)\) can you factor \(2x^2 -2x-1\) ?

OpenStudy (mymakersdaughter):

can you factor the 2 out?

hartnn (hartnn):

sure, we can.

hartnn (hartnn):

\(\color{blue}{\text{Originally Posted by}}\) @hartnn yep, notice you could factor out the 2 as well. \(2x(2x^2 -x-1)\) can you factor \(2x^2 -2x-1\) ? \(\color{blue}{\text{End of Quote}}\) i meant can you factor \(2x^2 -x-1\) ?

OpenStudy (mymakersdaughter):

I don't know how we could factor that any more because nothing can go into the 1

hartnn (hartnn):

can you find 2 numbers whose sum is -1 and product is -2 ? hint : only one of them is negative.

OpenStudy (mymakersdaughter):

-2 and 1?

OpenStudy (mymakersdaughter):

or -1 and 2?

hartnn (hartnn):

-2 and 1 is correct! so lets do factoring, \(2x^2 -2x -x-1 \\ 2x(x+1) -1(x+1)\) got that? can you proceed?

OpenStudy (mymakersdaughter):

so is that how we get the -2 as a root? How do we get the multiplicity?

hartnn (hartnn):

-2 is not the root! 2x(x+1) -1 (x+1) = 0 gives, (2x-1)(x+1) = 0 which means 2x-1 = 0 or x+1 = 0 so now can you find 2 other values of x ? (we already had x=0 before)

OpenStudy (mymakersdaughter):

Oh okay so we plug 0 into all of the x places?

hartnn (hartnn):

we get x from 2x-1 = 0 so, 2x = 1 x =1/2 x+1 = 0 so, x = -1 so we have 3 roots of x as -1, 0 , 1/2 got this?

OpenStudy (mymakersdaughter):

yes i've got it

hartnn (hartnn):

now about the mulitplicity, it is how many times that particular root repeats. for example, in 4,4,7. The multiplicity of 4 = 2, because it repeats 2 times. whereas the multiplicity of 7 is just 1.

OpenStudy (mymakersdaughter):

Ohhhhh!! okay so because no number repeats it has a multiplicity of 1?

hartnn (hartnn):

yup! each root has the multiplicity of 1.

OpenStudy (mymakersdaughter):

Yay!!!! Thank you soooo much for your help!!!! :D <3

hartnn (hartnn):

welcome ^_^